# Predictive Coherence and the Moment Hierarchy: Martingale Posteriors for Exchangeable Bernoulli Sequences

### 05 Mar 2026 v2

@@ 7.1 Complete class theorem
Theorem 7.2 (Blackwell–Girshick).
@@
Part (a) is standard. Part (b) follows from Blackwell–Girshick (1954) (compact parameter space, continuous risk). Uniqueness follows from strict propriety (Gneiting and Raftery, 2007). ∎
@@ 7.2 Plug-in dominance
Proposition 7.3 (Plug-in is strictly dominated).
@@
Fix $X_{1:n}$ with $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$ (which holds $\mathbb{P}_{\theta}$-a.s. since $\Pi$ has full support). The Bayes predictive is $\hat{P}^{\mathrm{Bayes}}(x)=\mathbb{E}[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}]$; the plug-in is $\hat{P}^{\mathrm{plug}}(x)=\theta_{n}^{s}(1-\theta_{n})^{k-s}$. By Corollary 6.5, for $k\geq 2$ these differ on a set of positive probability whenever the posterior is non-degenerate. Since the Bayes predictive equals the conditional distribution of the future block given $\mathcal{F}_{n}$, strict propriety implies that $\hat{P}^{\mathrm{Bayes}}$ strictly minimises expected score. Hence the plug-in’s expected score is strictly higher. Integrating over $\mathcal{F}_{n}$ under $\mathbb{P}_{\theta}$ preserves the strict inequality. ∎
The dominance is strict whenever $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$. For a geometric interpretation of admissibility via risk-set boundaries and supporting hyperplanes, and its extension to sequential, coverage, and approachability criteria, see Polson and Zantedeschi (2026).
Corollary 7.4 (Inadmissibility).
@@
Immediate from Proposition 7.3: the Bayes rule strictly dominates the plug-in whenever $\Pi$ has full support. ∎
Corollary 7.5 (Mean-only rules).
@@
By Theorem 6.3, two posteriors sharing the same mean generically yield different Bayes-optimal predictives for $k\geq 2$. A mean-only rule assigns the same predictive to both; it cannot be Bayes-optimal for both simultaneously. ∎
Remark 7.6 (Log-score risk gap).
At $k=2$, the Bayes predictive assigns run probability $p_{B}=\mathbb{E}[(1-\theta)^{2}\mid\mathcal{F}_{n}]=(1-m_{n})^{2}+\sigma_{n}^{2}$ and the plug-in assigns $p_{P}=(1-m_{n})^{2}$. The KL divergence between these Bernoulli predictives satisfies $D_{\mathrm{KL}}(\text{Bayes}\|\text{plug-in})\geq\sigma_{n}^{4}/(2\ln 2)$ by Pinsker’s inequality, giving a lower bound on the risk gap that is quadratic in the posterior variance.
Example 7.7 (Numerical risk gap).
@@
This KL divergence, while small in absolute terms, is strictly positive and cannot be eliminated by any mean-only adjustment. At $k=4$ the divergence grows to approximately $0.021$, an order of magnitude larger. Accumulated over a sequential prediction task with $T$ rounds, the total excess log-loss is $\Omega(T\sigma_{n}^{4})$, non-negligible when $n$ is moderate relative to $T$.
Remark 7.8 (Brier score risk gap).
@@
Under the log score, the loss from using $q$ instead of $p$ equals $D_{\mathrm{KL}}(p\|q)$ (Cover and Thomas, 2006): KL is simultaneously predictive regret and the canonical divergence between distributions. The Bayes predictive minimizes expected loss, and the gap to any plug-in baseline is a KL term (Proposition 7.3, Remark 7.6).
@@ 8 Hill’s $A_{(n)}$ as a positive example
Proposition 8.1 (Hill/Jeffreys one-step predictive).
@@
The posterior mean of $\mathrm{Beta}(a,b)$ is $a/(a+b)$. With $a=\tfrac{1}{2}+S_{n}$ and $b=\tfrac{1}{2}+n-S_{n}$, the result follows by Theorem 4.1 with $k=1$. ∎
Proposition 8.2 ($k$-step run probabilities).
@@
By Theorem 4.1 and the Beta moment formula: $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]=B(a_{n},b_{n}+k)/B(a_{n},b_{n})=(b_{n})_{k}/(N_{n})_{k}$. ∎
@@ 8.1 Full moment recovery
Hill’s rule specifies every conditional moment as an explicit function of $(n,S_{n})$. It is a Bayes rule and therefore admissible (Theorem 7.2(a)).
@@ 8.2 Plug-in versus Bayes comparison
Remark 8.3 (Lane–Sudderth caveat).
@@
Remark 8.4 (Red/black card prediction).
In the standard deck-prediction problem, one-step prediction depends only on the posterior mean (martingale coherence), but $k$-step run probabilities require $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$, hence all posterior moments up to order $k$ (Corollary 4.2). The curvature of $(1-\theta)^{k}$ makes the variance contribution non-negligible whenever the remaining composition deviates from $1/2$.
@@ 9 Asymptotic reconciliation
The moment insufficiency of Section 6 is a finite-sample phenomenon.
Proposition 9.1 (Asymptotic agreement).
@@
By Proposition 6.4. ∎
@@
Proposition 9.2 (Uniform bound).
@@
By Proposition 6.4, $\sup_{k\leq K_{n}}\frac{k(k-1)}{2}\sigma_{n}^{2}\leq\frac{K_{n}^{2}}{2}\cdot O(1/n)\to 0$ when $K_{n}=o(\sqrt{n})$. ∎
Proposition 9.3 (Bernstein–von Mises rate).
Let $\Pi$ have a continuous, positive density near $\theta_{0}\in(0,1)$. For fixed $k$,
$\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}=\frac{k(k-1)}{2}(1-\theta_{0})^{k-2}\cdot\frac{\theta_{0}(1-\theta_{0})}{n}+O(n^{-2})\quad\text{a.s.}$
The relative discrepancy is $O(1/n)$.
@@
By the Bernstein–von Mises theorem, $\sigma_{n}^{2}=\theta_{0}(1-\theta_{0})/n+O(n^{-2})$ a.s. Substituting into Proposition 6.4 and noting $\xi_{n}\to\theta_{0}$ a.s. gives the expansion. ∎
Remark 9.4 (Finite-sample relevance).
The discrepancy vanishes as $n\to\infty$ for fixed $k$. When the prediction horizon grows with $n$, Proposition 9.2 requires $k=o(\sqrt{n})$; for $k\geq C\sqrt{n}$, the bound $\frac{k(k-1)}{2}\sigma_{n}^{2}$ remains $O(1)$.
@@ 10 Structural hierarchy
Figure 2: The predictive moment hierarchy. Bottom: the martingale / mean-only tier (cf. linear Bayes) constrains only the first moment and determines one-step predictives. Middle: specifying $J$ moments (Goldstein (1975a) prevision tier) determines $k$-step predictives for $k\leq J$ but not for $k>J$ (Theorem 6.3). Top: the full conditional law determines all $k$-step predictives (predictive completeness; Theorem 10.3).
@@ 10.1 Hierarchy of predictive frameworks
The hierarchy is strict: specifying $j$ moments determines $k$-step predictives for $k\leq j$ (Corollary 4.2) but not for $k>j$ (Theorem 6.3 at order $j+1$). The middle tier corresponds to linear Bayes / prevision updating (cf. Goldstein (1975a)).
Remark 10.1 (Conformal row).
@@ 10.2 Closure theorem
Definition 10.2 (Predictive completeness).
@@
Theorem 10.3 (Closure).
Let $(\theta_{n})$ be a $[0,1]$-valued martingale satisfying Assumption 3.4, with terminal value $\theta_{\infty}$. The following are equivalent:
@@
A martingale posterior is predictively complete if and only if it uniquely determines the conditional law of $\theta_{\infty}$ given $\mathcal{F}_{n}$. Condition (1) is necessary but not sufficient: it fixes only $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ and leaves the conditional law underdetermined. On $[0,1]$, specifying the law is equivalent to specifying all moments (Hausdorff determinacy; Theorem A.1), which is equivalent to specifying all $k$-step predictives.

### 01 Apr 2026 v5

@@ 7.1 Complete class theorem
Theorem 29 (Blackwell–Girshick).
@@
Part (a) follows by integrating against $\Pi$: if $\delta^{\prime}$ dominated $\delta^{*}$, it would have lower Bayes risk, contradicting optimality. Part (b) follows from compactness of $[0,1]$ and continuity of risk; see Blackwell and Girshick (1954). The Bayes-optimal predictive is the unique minimiser of expected score under strict propriety, which by Theorem 13 equals $\mathbb{E}[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}]$. ∎
@@ 7.2 Plug-in dominance
Proposition 30 (Plug-in is strictly dominated).
@@
Fix $X_{1:n}$ with $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$ (which holds $\mathbb{P}_{\theta}$-a.s. since $\Pi$ has full support). The Bayes predictive is $\hat{P}^{\mathrm{Bayes}}(x)=\mathbb{E}[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}]$; the plug-in is $\hat{P}^{\mathrm{plug}}(x)=\theta_{n}^{s}(1-\theta_{n})^{k-s}$. By Corollary 25, for $k\geq 2$ these differ on a set of positive probability whenever the posterior is non-degenerate. Since the Bayes predictive equals the conditional distribution of the future block given $\mathcal{F}_{n}$, strict propriety implies that $\hat{P}^{\mathrm{Bayes}}$ strictly minimises expected score. Hence the plug-in’s expected score is strictly higher. Integrating over $\mathcal{F}_{n}$ under $\mathbb{P}_{\theta}$ preserves the strict inequality. ∎
The dominance is strict whenever $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$.
Corollary 31 (Inadmissibility).
@@
By Proposition 30, for every $\theta\in(0,1)$, the Bayes predictive achieves strictly lower risk than the plug-in predictive under any strictly proper scoring rule. Since the Bayes rule uniformly dominates the plug-in and is itself admissible (Theorem 29(a)), the plug-in is inadmissible. ∎
Corollary 32 (Mean-only rules).
@@
By Theorem 23, there exist posteriors $\nu_{1},\nu_{2}$ with $\int\theta\,\mathrm{d}\nu_{1}=\int\theta\,\mathrm{d}\nu_{2}=m_{n}$ but $\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}\neq\int(1-\theta)^{k}\,\mathrm{d}\nu_{2}$ for $k\geq 2$. A mean-only rule $\delta$ maps both posteriors to the same predictive $\hat{P}_{\delta}(m_{n})$. By strict propriety, the unique Bayes-optimal predictive under $\nu_{i}$ is $\hat{P}^{*}_{i}(x)=\int\theta^{s}(1-\theta)^{k-s}\,\mathrm{d}\nu_{i}$, and $\hat{P}^{*}_{1}\neq\hat{P}^{*}_{2}$. Since $\delta$ assigns the same $\hat{P}$ to both, it cannot equal both $\hat{P}^{*}_{1}$ and $\hat{P}^{*}_{2}$; hence $\delta$ is suboptimal under at least one posterior and is outside the Bayes class (unless $\nu_{1}$ or $\nu_{2}$ is degenerate). ∎
Remark 33 (Log-score risk gap).
At $k=2$, the Bayes predictive assigns run probability $p_{B}=\mathbb{E}[(1-\theta)^{2}\mid\mathcal{F}_{n}]=(1-m_{n})^{2}+\sigma_{n}^{2}$ and the plug-in assigns $p_{P}=(1-m_{n})^{2}$. Since $p_{B}\neq p_{P}$ whenever $\sigma_{n}^{2}>0$, the KL divergence $D_{\mathrm{KL}}(\text{Bayes}\|\text{plug-in})$ is strictly positive, so the log-score risk gap is strictly positive whenever the posterior is non-degenerate.
Example 34 (Numerical risk gap).
@@
This KL divergence, while small in absolute terms, is strictly positive and cannot be eliminated by any mean-only adjustment. At $k=4$ the divergence grows to approximately $0.021$, an order of magnitude larger.
Remark 35 (Brier score risk gap).
@@
Under the log score, the plug-in–Bayes gap equals $D_{\mathrm{KL}}(\text{Bayes}\|\text{plug-in})$ (Cover and Thomas, 2006), connecting predictive regret to KL divergence.
Figure 3: Scoring-rule regret (plug-in minus Bayes). Top left: KL regret $D_{\mathrm{KL}}(\hat{P}^{\mathrm{Bayes}}\|\hat{P}^{\mathrm{plug}})$ vs. horizon $k$ at $S_{n}/n\approx 0.4$ under the Jeffreys prior, for sample sizes $n\in\{5,10,20,50,100\}$. Regret grows with $k$ (more moments neglected) and shrinks with $n$ (posterior concentrates). Top right: Brier regret $(p_{B}-p_{P})^{2}$, same cases. Bottom row: regret vs. posterior variance across all prior/sample-size combinations, showing the $O(\sigma_{n}^{4})$ Brier scaling (Remark 35) and the strictly positive KL gap whenever $\sigma_{n}^{2}>0$ (Remark 33). All values are exact.
@@ 8 Hill’s $A_{(n)}$ as a positive example
Proposition 36 (Hill/Jeffreys one-step predictive).
@@
The posterior mean of $\mathrm{Beta}(a,b)$ is $a/(a+b)$. With $a=\tfrac{1}{2}+S_{n}$ and $b=\tfrac{1}{2}+n-S_{n}$, the result follows by Theorem 13 with $k=1$. ∎
Proposition 37 ($k$-step run probabilities).
@@
By Theorem 13 and the Beta moment formula: $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]=B(a_{n},b_{n}+k)/B(a_{n},b_{n})=(b_{n})_{k}/(N_{n})_{k}$. ∎
@@ 8.1 Full moment recovery
Hill’s rule specifies every conditional moment as an explicit function of $(n,S_{n})$. It is a Bayes rule and therefore admissible (Theorem 29(a)).
@@ 8.2 Plug-in versus Bayes comparison
Figure 4: Bayes vs. plug-in predictive gap under Beta-Bernoulli updating. Left: absolute gap $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}$ vs. horizon $k$ for the Jeffreys prior at $S_{n}/n\approx 0.4$, extending Table 2 across sample sizes $n\in\{5,10,20,50,100\}$. The gap grows with $k$ (each step adds a new undetermined moment) and shrinks with $n$ (posterior concentrates). Centre: the corresponding relative gap, which exceeds $50\%$ at $k=8$ for $n=5$, illustrating that the plug-in can substantially underestimate the Bayes predictive at moderate horizons. Right: absolute gap vs. posterior variance across all priors and configurations, confirming the $O(\sigma_{n}^{2})$ scaling from Proposition 24. All values are computed exactly via $(b_{n})_{k}/(N_{n})_{k}$.
Remark 38 (Lane–Sudderth caveat).
@@
Remark 39 (Red/black card prediction).
In the standard deck-prediction problem, one-step prediction depends only on the posterior mean (martingale coherence), but $k$-step run probabilities require $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$, hence all posterior moments up to order $k$ (Corollary 14). The curvature of $(1-\theta)^{k}$ makes the variance contribution non-negligible whenever the remaining composition deviates from $1/2$.
@@ 9 Asymptotic reconciliation
The moment insufficiency of Sections 67 is a finite-sample structural phenomenon. It does not preclude asymptotic proximity between empirical and predictive measures, which has been studied extensively.
Berti, Crimaldi, Pratelli, and Rigo (2009) study the empirical process $C_{n}(B)=\sqrt{n}\{\mu_{n}(B)-\mathbb{P}(X_{n+1}\in B\mid\mathcal{F}_{n})\}$ for c.i.d. and exchangeable sequences, obtaining stable convergence at rate $n^{-1/2}$. Berti, Pratelli, and Rigo (2004) establish law-level rates using bounded Lipschitz metrics, and Berti, Pratelli and Rigo (2017) give matching bounds for finite alphabets: in the Bernoulli case,
$\frac{a}{n}\;\leq\;\rho\bigl[\mathcal{L}(\mu_{n}),\,\mathcal{L}(\mu)\bigr]\;\leq\;\frac{b}{n+1},$ (20)
where $a=\mathbb{E}[\theta(1-\theta)]$ and $b$ depends on the total variation of the prior density. The lower bound is driven by the expected posterior variance, the same quantity that governs the moment-hierarchy discrepancy in Proposition 24. See also Berti, Pratelli and Rigo (2018) and Leisen, Pratelli and Rigo (2026) for further asymptotic developments.
Our concern is complementary: we ask what finite-horizon predictive content is identified by first-moment information itself, independent of any asymptotic limit. The bounds of (2009) and (2017) measure how fast empirical and predictive measures approach one another; our results identify which functionals of the posterior (moments of order $\geq 2$) remain structurally underdetermined by the first moment alone. Despite this structural incompleteness, posterior concentration brings plug-in and Bayes predictives close together asymptotically, though this closeness does not erase the hierarchy.
Proposition 40 (Asymptotic agreement).
@@
By Proposition 24 with $f(\theta)=(1-\theta)^{k}$, the discrepancy satisfies $|\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}|=\frac{k(k-1)}{2}(1-\xi_{n})^{k-2}\sigma_{n}^{2}$ for some $\xi_{n}\in[0,1]$. Since $(1-\xi_{n})^{k-2}\leq 1$, the bound $\frac{k(k-1)}{2}\sigma_{n}^{2}$ holds. As $\sigma_{n}^{2}\to 0$ a.s. by assumption, the right-hand side vanishes for each fixed $k$, giving the claimed convergence. ∎
@@
Proposition 41 (Uniform bound).
@@
By Proposition 24, for each $k\leq K_{n}$, $|\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}|\leq\frac{k(k-1)}{2}\sigma_{n}^{2}$. Taking the supremum over $1\leq k\leq K_{n}$ replaces k(k1)/2k(k-1)/2 by its maximum $K_{n}(K_{n}-1)/2\leq K_{n}^{2}/2$, giving $\sup_{k\leq K_{n}}|\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}|\leq\frac{K_{n}^{2}}{2}\sigma_{n}^{2}$. Under $\mathrm{Beta}(a,b)$ (or any prior with $\sigma_{n}^{2}=O(1/n)$), the right-hand side is $O(K_{n}^{2}/n)$, which converges to zero if and only if $K_{n}=o(\sqrt{n})$. This threshold reflects the tension between growing curvature in $(1-\theta)^{k}$ and decreasing posterior variance. ∎
Proposition 42 (Bernstein–von Mises rate).
Let $\theta_{0}\in(0,1)$ and suppose that the prior $\Pi$ admits a continuous, strictly positive density in a neighbourhood of $\theta_{0}$. Under $\mathbb{P}_{\theta_{0}}$ (that is, when data are generated i.i.d. $\mathrm{Bern}(\theta_{0})$ and posterior inference uses the prior $\Pi$), for fixed $k\geq 2$ and as $n\to\infty$,
$\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}=O(n^{-1})\quad\mathbb{P}_{\theta_{0}}\text{-a.s.}$
The discrepancy between the Bayes predictive and the plug-in predictive thus vanishes at rate $O(1/n)$.
@@
Under $\mathbb{P}_{\theta_{0}}$, the observations are i.i.d. $\mathrm{Bern}(\theta_{0})$, and the classical Bernstein–von Mises theorem (van der Vaart, 1998, Chapter 10) gives the posterior variance $\sigma_{n}^{2}=O(n^{-1})$ $\mathbb{P}_{\theta_{0}}$-a.s.; the regularity conditions (continuous, strictly positive density near $\theta_{0}$) are assumed. By Proposition 24,
$0\;\leq\;\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}\;\leq\;\frac{k(k-1)}{2}\,\sigma_{n}^{2}\;=\;O(n^{-1}).$
The lower bound is strict when $\sigma_{n}^{2}>0$ (Proposition 24), and the upper bound uses $(1-\xi_{n})^{k-2}\leq 1$. ∎
Remark 43 (Role of $\theta_{0}$).
The $O(n^{-1})$ rate in Proposition 42 holds under $\mathbb{P}_{\theta_{0}}$ for each fixed $\theta_{0}\in(0,1)$. The implicit constant depends on $\theta_{0}$ (through $\sigma_{n}^{2}\sim\theta_{0}(1-\theta_{0})/n$) and on $k$. If $\Pi$ has full support, the regularity conditions hold for every $\theta_{0}\in(0,1)$ simultaneously; there is no contradiction, because for each $\theta_{0}$ the convergence is under the corresponding product measure $\mathbb{P}_{\theta_{0}}$, which generates different data sequences and different posterior sequences. The left-hand side is a computable posterior functional that does not require knowledge of $\theta_{0}$, but its $\mathbb{P}_{\theta_{0}}$-a.s. rate depends on the true parameter through the data it generates.
Remark 44 (Finite-sample relevance).
The discrepancy vanishes as $n\to\infty$ for fixed $k$. When the prediction horizon grows with $n$, Proposition 41 requires $k=o(\sqrt{n})$; for $k\geq C\sqrt{n}$, the bound $\frac{k(k-1)}{2}\sigma_{n}^{2}$ remains $O(1)$. In short, the plug-in and Bayes predictives agree uniformly over all horizons that grow slower than $\sqrt{n}$, but diverge once the horizon reaches the $\sqrt{n}$ scale: first-moment information suffices for short-range prediction but not for long-range prediction, even asymptotically.
Figure 5: Asymptotic reconciliation under Jeffreys prior, averaged over 200 simulated Bernoulli sequences. Top row: discrepancy $\Delta_{n,k}$ vs. sample size $n$ for fixed horizons $k\in\{2,4,8\}$ at three true parameter values $\theta_{0}\in\{0.3,0.5,0.7\}$. The discrepancy vanishes at rate $O(1/n)$ (Proposition 42), with larger $k$ producing a proportionally larger gap at each $n$. Bottom row: growing horizon $k_{n}\approx c\sqrt{n}$ for $c\in\{0.5,1.0,2.0\}$. When $k_{n}=o(\sqrt{n})$ the discrepancy still vanishes; at $k_{n}\sim c\sqrt{n}$ it stabilises at a positive level, illustrating the $\sqrt{n}$ threshold of Proposition 41: first-moment information suffices for short-range prediction but not for long-range prediction.
@@ 10 Structural hierarchy
Figure 6: The predictive moment hierarchy. Bottom: the martingale / mean-only tier (cf. linear Bayes) constrains only the first moment and determines one-step predictives. Middle: specifying $J$ moments (Goldstein (1975a) prevision tier) determines $k$-step predictives for $k\leq J$ but not for $k>J$ (Theorem 23). Top: the full conditional law determines all $k$-step predictives (predictive completeness; Theorem 47).
@@ 10.1 Hierarchy of predictive frameworks
The hierarchy is strict: specifying $j$ moments determines $k$-step predictives for $k\leq j$ (Corollary 14) but not for $k>j$ (Theorem 23 at order $j+1$). The middle tier corresponds to linear Bayes / prevision updating (cf. Goldstein (1975a)).
Remark 45 (Conformal row).
@@ 10.2 Closure theorem
Definition 46 (Predictive completeness).
@@
Theorem 47 (Closure).
Let $(\theta_{n})$ be a $[0,1]$-valued martingale satisfying Assumption 5, with terminal value $\theta_{\infty}$. The following are equivalent:
@@
A martingale posterior is predictively complete if and only if it uniquely determines the conditional law of $\theta_{\infty}$ given $\mathcal{F}_{n}$. Condition (1) is necessary but not sufficient: it fixes only $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ and leaves the conditional law underdetermined. On $[0,1]$, specifying the law is equivalent to specifying all moments (Hausdorff determinacy; Theorem 51), which is equivalent to specifying all $k$-step predictives.